A kid's toy store found that its quarterly sales of $1,000 can be adequately forecasted by the following multiplicative seasonal model.

Trend Component: T = 150 + 35t

Seasonal Index: Q1= .81 Q2= 1.09 Q3= .91 Q4= 1.19

Origin = Q4, 2012

Units (t) = Quarterly

A. Forecast sales for 2019
Answer #1

From 2012 to 2018, 6 years have elapsed, which means 6*4 = 24 quarters.

So time index (t) for the 4 quarters of year 2019 are 25 to 28

Quarter 1, 2019 forecast = (150 + 35*25)*0.81 = 830.25

Quarter 2, 2019 forecast = (150 + 35*26)*1.09 = 1155.40

Quarter 3, 2019 forecast = (150 + 35*27)*0.91 = 996.45

Quarter 4, 2019 forecast = (150 + 35*28)*1.19 = 1344.70

Total sales forecast for year 2019 = 830.25 + 1155.40 + 996.45 + 1344.70 = 4326.80